### v1

@@ Abstract
The microscopic composition of neutron star interiors remains uncertain, with viable possibilities ranging from purely nucleonic matter to hyperonic matter, dark‑matter‑admixed cores, and strange self‑bound matter. Traditional constraints on the equation of state (EOS) rely on Tolman–Oppenheimer–Volkoff (TOV) modelling and comparison with multimessenger observations, but machine learning (ML) provides a complementary pathway by learning composition‑dependent patterns directly from astrophysically accessible observables.
This work introduces a compact supervised‑learning framework for EOS‑family classification using stellar properties derived from TOV modelling, asteroseismology, and gravitational‑wave descriptors. A labelled dataset of neutron star configurations spanning four EOS families: nucleonic, strange‑matter, dark‑matter‑admixed, and hyperonic, is constructed using seven input features: gravitational mass $M$, radius $R$, fundamental $f$-mode frequency, quadrupole moment $Q$, redshift $z$, damping time $\tau$, and characteristic strain $h_{c}$. A multilayer perceptron is trained to infer the underlying matter composition.
On a held‑out test set, the classifier attains an accuracy of 97.4% with strong class‑wise precision and recall. Permutation‑based feature‑importance analysis shows that oscillation‑related quantities, especially the $f$-mode frequency and damping time—dominate the discriminatory power, while mass and radius provide secondary support. Residual misclassifications cluster in physically intuitive regions where different EOS families yield partially overlapping macroscopic signatures.
The results demonstrate that even lightweight neural models can reliably identify EOS‑family fingerprints from a modest set of observables, offering a reproducible baseline for future extensions incorporating Bayesian uncertainty and observational posteriors from NICER and gravitational‑wave events.
@@ 1 Introduction
Each scenario imprints characteristic signatures on macroscopic observables such as the mass-radius ($M$-$R$) relation and the f-mode frequency Andersson and Kokkotas, (1998); Doneva et al., (2015); Kokkotas and Schmidt, (1999) , which can be constrained by X-ray timing, radio pulsar measurements , and gravitational waves from binary mergers.
Traditional studies proceed by specifying an equation of state (EOS), solving the Tolman-Oppenheimer-Volkoff (TOV) equations for hydrostatic equilibrium, and comparing the resulting $M$-$R$ curves with observational constraints. In recent years, machine learning (ML) techniques have been deployed in neutron star physics Fujimoto et al., (2018); Morawski and Bejger, (2020); Raithel et al., (2017); Krastev, (2022); Soma et al., (2023) , targeting tasks such as EOS reconstruction from mock observations Fujimoto et al., (2021), probabilistic inference of the speed-of-sound profile, or mapping between observational posteriors and internal structure parameters. These approaches demonstrate that ML can encapsulate complex non-linear relationships between the EOS and observables, but often rely on high-dimensional EOS parametrisations or large training sets.
@@
  • Constructed a labelled dataset of neutron-star models spanning four EOS families (nucleonic, hyperonic, non-self-interactive dark-matter-admixed, and strange matter), each characterised by seven observables: $(M,R,f\text{-mode},Q,\text{redshift},\tau,h_{c})$.
@@ 2.2 TOV equations and mass-radius relations
For each stable stellar model we store the EOS-family label together with the set of observables used for classification. In the present dataset these include the gravitational mass $M$, radius $R$, the fundamental $f$-mode frequency, quadrupole moment, redshift, the $f$-mode damping time $\tau$, and the characteristic strain $h_{c}$.
Figure 1: Representative mass-radius relations for the four EOS families used to construct the machine-learning dataset. While the curves partially overlap in certain mass ranges, composition-dependent trends persist, particularly at intermediate and high masses, motivating the supervised classification approach adopted in this work.
Figure 1 shows representative mass-radius relations for the four EOS families considered in this work. As expected, different compositions give rise to systematically different stiffness and compactness trends, with nucleonic models typically producing larger radii at a given mass, while hyperonic and strange-matter EOSs exhibit softer behaviour at high densities. Dark-matter-admixed configurations can mimic either stiff or soft baryonic EOSs depending on the dark-sector parameters, leading to partial overlap with the other families. Importantly, the mass-radius curves are not fully disjoint across the astrophysically relevant mass range. In particular, multiple EOS families can support neutron stars with similar masses and radii near $M\sim 1.4$-$2.0\,M_{\odot}$. This overlap demonstrates that classification based on mass and radius alone is intrinsically ambiguous and motivates the inclusion of additional observables, such as oscillation-related quantities, in the machine-learning analysis. The supervised classification task addressed in this work is therefore non-trivial and probes genuinely composition-dependent information beyond simple mass-radius trends.
@@ 3.1 Dataset and features
The dataset consists of theoretically generated neutron-star models spanning four EOS families: nucleonic matter, strange matter, dark-matter–admixed matter, and hyperonic matter. For each EOS family, we generate approximately $190$ stable neutron-star models by sampling the central energy density over a range of values supporting a $\sim 1\,M_{\odot}$ star up to the maximum-mass configuration of the EOS. This results in a total dataset size of $\sim 770$ samples across all four EOS families. Only configurations on the stable branch of the mass–central-density relation are retained. For each stellar configuration, we extract seven physically motivated observables: gravitational mass, stellar radius, fundamental $f$-mode frequency, a quadrupole moment, redshift, the $f$-mode damping time ($\tau$), and the characteristic gravitational-wave strain.
Prior to training, samples with missing feature values are removed. All features are standardised to zero mean and unit variance using statistics computed from the training set only, which improves numerical conditioning and prevents information leakage.
3.2 Train-validation-test split
@@ 3.3 Model architecture and training
A fully connected multilayer perceptron (MLP) classifier is implemented in scikit-learn. The network consists of two hidden layers with 64 neurons each and ReLU activations. Training uses the Adam optimiser with early stopping based on validation loss to reduce overfitting. The model outputs class probabilities over the four EOS families.
@@ 4 Results
This section reports the performance of the supervised machine-learning classifier trained to identify the underlying neutron-star matter composition from macroscopic and oscillation-related observables. The input features include the gravitational mass, stellar radius, fundamental $f$-mode frequency, quadrupole moment, redshift, damping time, and characteristic gravitational-wave strain. The output classes correspond to four equation-of-state (EOS) families: nucleonic matter, strange matter, dark-matter–admixed matter, and hyperonic matter.
@@ 4.2 Overall classification performance
The classifier achieves a test-set accuracy of approximately $97\%$, demonstrating that the chosen observables encode strong information about the underlying matter composition. In the present configuration, the held-out test accuracy is $97.4\%$, evaluated on a stratified test set containing $N_{\text{test}}=116$ samples. The detailed class-wise performance is summarised by the confusion matrix shown in Fig. 3.
@@ 4.6 Feature importance
These results indicate that while global structural properties such as mass and radius contain useful information, they are insufficient on their own to robustly distinguish between EOS families. Oscillation-related quantities encode additional composition-dependent physics and play a crucial role in breaking degeneracies that arise when different EOSs produce similar mass–radius relations. Physically, the enhanced importance of oscillation-related observables can be understood from their sensitivity to the stellar compactness and internal sound-speed profile, both of which depend directly on the underlying EOS. While mass and radius provide integral constraints on stellar structure, mode frequencies and damping times probe the dynamical response of the star and are therefore more sensitive to composition-dependent changes in dense matter. This highlights the potential of neutron-star asteroseismology as a discriminator between competing EOS scenarios.
@@ 5 Discussion and Conclusions
The present work intentionally provides a minimal, reproducible baseline rather than a full end-to-end observational inference engine. Straightforward extensions include: (i) training on samples drawn from observational posteriors (NICER and GW events) Miller et al., (2019); Abbott et al., (2018) rather than point estimates; (ii) uncertainty-aware models (e.g., Bayesian neural networks) to calibrate predictive confidence; and (iii) inclusion of additional observables (e.g., higher-order Love numbers or multi-mode asteroseismology) to break degeneracies between hyperonic and strange-matter scenarios. These results support the role of machine learning as a complementary tool that can accelerate EOS-family inference and highlight regions where multimessenger data are most informative.

### v2

@@ Abstract
We investigate the extent to which supervised machine-learning techniques can distinguish between neutron-star matter models using macroscopic and oscillation-related quantities derived from theoretical stellar configurations. Four representative matter scenarios nucleonic, hyperonic, dark-matter–admixed, and strange-matter models are considered, and a synthetic dataset is constructed from solutions of the Tolman–Oppenheimer–Volkoff equations under fixed microphysical and transport assumptions.
A shallow neural-network classifier is trained on physically motivated features, including gravitational mass, stellar radius, and oscillation-related quantities, to evaluate classification performance across the model space. Rather than aiming at unique composition inference, the analysis focuses on identifying regimes of distinguishability and intrinsic degeneracy between models. We find that certain matter scenarios can be separated under controlled assumptions, while others exhibit substantial overlap, reflecting fundamental similarities in their effective equations of state.
These results demonstrate that machine learning provides a useful computational framework for mapping the limits of model classification in neutron-star studies, clarifying where inference is feasible and where it remains intrinsically model dependent. The methodology is readily extensible to more complex microphysics and to future multi-messenger datasets.
@@ 1 Introduction
Each scenario imprints characteristic signatures on macroscopic observables such as the mass-radius ($M$-$R$) relation and the f-mode frequency (Andersson and Kokkotas, (1998); Doneva et al., (2015); Kokkotas and Schmidt, (1999)), which can be constrained by X-ray timing, radio pulsar measurements, and gravitational waves from binary mergers.
Traditional studies proceed by specifying an equation of state (EOS), solving the Tolman-Oppenheimer-Volkoff (TOV) equations for hydrostatic equilibrium, and comparing the resulting $M$-$R$ curves with observational constraints. In recent years, machine learning (ML) techniques have been deployed in neutron star physics (Fujimoto et al., (2018); Morawski and Bejger, (2020); Raithel et al., (2017); Krastev, (2022); Soma et al., (2023)) , targeting tasks such as EOS reconstruction from mock observations Fujimoto et al., (2021), probabilistic inference of the speed-of-sound profile, or mapping between observational posteriors and internal structure parameters. These approaches demonstrate that ML can encapsulate complex non-linear relationships between the EOS and observables, but often rely on high-dimensional EOS parametrisations or large training sets.
@@
  • Constructed a labelled dataset of neutron-star models spanning four EOS families (nucleonic, hyperonic, non-self-interactive dark-matter-admixed, and strange matter), each characterised by seven observables: $(M,R,f\text{-mode},$ mode quadrupole coefficient $Q,\text{redshift},$ damping time $\tau,$ $h_{c})$. The inclusion of oscillation-based quantities extends beyond the purely structural observables typically used in previous ML studies. A deliberately minimal multilayer perceptron architecture is employed to isolate physically driven separability rather than architecture-dependent performance gains.
@@ 2 Equations of State and Stellar Models
In this work, we distinguish between observables primarily controlled by the equation of state (e.g., mass, radius, and oscillation frequencies) and quantities that can additionally depend on transport assumptions (e.g., damping times). Transport parameters are held fixed across all EOS families considered here. Consequently, variations in oscillation-related quantities reflect structural differences induced by the EOS within this controlled modelling framework. The classifier is therefore designed to learn relative patterns within a fixed theoretical setup rather than provide EOS-independent or observationally unique inference.
@@ 2.2 TOV equations and mass-radius relations
For each stable stellar model we store the EOS-family label together with the set of observables used for classification.
In the present dataset these include the gravitational mass $M$, radius $R$, the fundamental $f$-mode frequency, mode quadrupole coefficient $Q$, redshift, the $f$-mode damping time $\tau$, and the characteristic strain $h_{c}$. The oscillation-related quantities ($f$-mode frequency, damping time $\tau$, quadrupole amplitude proxy $Q$, and characteristic strain $h_{c}$) are computed using the same perturbative framework and modelling assumptions presented in Ref. Husain, (2026), where detailed derivations and numerical procedures are provided. In particular, transport prescriptions and damping mechanisms are held fixed across all EOS families considered here. The present work does not introduce new oscillation calculations; rather, it reuses these consistently computed quantities to construct a supervised classification dataset. Consequently, differences between EOS families reflect structural variations within the adopted theoretical framework rather than variations in microphysical transport parameters. The quadrupole quantity $Q$ refers to the dimensionless mode-normalisation coefficient entering the gravitational-wave luminosity and strain expressions for nonradial oscillations. The quadrupole quantity used here arises from nonradial $f$-mode oscillations of a spherically symmetric background star and does not require stellar rotation. It represents the time-dependent mass quadrupole entering the gravitational-wave emission formalism. The characteristic strain $h_{c}$ is computed for a fiducial source distance following the prescription detailed in Husain, (2026), ensuring consistent treatment of oscillation amplitudes across all EOS families.
Figure 1: Representative mass-radius relations for the four EOS families used to construct the machine-learning dataset. Note that absolute radii are systematically reduced due to the absence of a low-density crust in the EOS construction Husain, (2026). This modelling choice follows previous work and does not affect relative trends between models, which are the quantities learned by the classifier.
Figure 1 shows representative mass-radius relations for the four EOS families considered in this work Husain, (2026). Because all EOS families are treated consistently without explicit crust matching, the absolute radii are systematically reduced relative to fully unified EOS models. However, this modelling choice is applied uniformly across all classes, and therefore relative structural differences between EOS families remain meaningful within the controlled classification framework adopted here. As expected, different compositions give rise to systematically different stiffness and compactness trends, with nucleonic models typically producing larger radii at a given mass, while hyperonic and strange-matter EOSs exhibit softer behaviour at high densities. Dark-matter-admixed configurations can mimic either stiff or soft baryonic EOSs depending on the dark-sector parameters, leading to partial overlap with the other families. Importantly, the mass-radius curves are not fully disjoint across the astrophysically relevant mass range. In particular, multiple EOS families can support neutron stars with similar masses and radii near $M\sim 1.4$-$2.0\,M_{\odot}$. This overlap demonstrates that classification based on mass and radius alone is intrinsically ambiguous and motivates the inclusion of additional observables, such as oscillation-related quantities, in the machine-learning analysis. The supervised classification task addressed in this work is therefore non-trivial and probes genuinely composition-dependent information beyond simple mass-radius trends.
@@ 3.1 Dataset and features
The dataset consists of theoretically generated neutron-star models spanning four EOS families: nucleonic, hyperonic, strange-matter, and dark-matter–admixed models.
For each EOS family, approximately $190$ stable stellar configurations are generated by sampling the central energy density over the range that supports stars from $\sim 1\,M_{\odot}$ up to the maximum stable mass. This results in a total dataset size of approximately $770$ samples across all classes. Only configurations on the stable branch of the mass–central-density relation are retained. Although the total dataset size is modest compared to typical deep-learning applications, the parameter space is low-dimensional and fully controlled. The present sample size is therefore sufficient to probe class separability and degeneracy structure within the adopted EOS families without requiring large-scale data augmentation.
For each stellar configuration, seven observables are extracted: gravitational mass $M$, stellar radius $R$, fundamental $f$-mode frequency, mode quadrupole coefficient $Q$, gravitational redshift, $f$-mode damping time $\tau$, and characteristic gravitational-wave strain $h_{c}$.
@@ 3.2 Model architecture and training
A fully connected multilayer perceptron (MLP) classifier is implemented in scikit-learn. The network consists of two hidden layers with 64 neurons each and ReLU activations. Training is performed using the Adam optimiser with an initial learning rate of $10^{-3}$ and L2 regularisation strength $\alpha=10^{-4}$. Mini-batch training is used with a batch size of 64.
A maximum of 500 training epochs is allowed, with early stopping enabled to mitigate overfitting. An internal validation fraction of 0.15 is used, and training is terminated if the validation loss does not improve for 20 consecutive epochs. All random splits and model initialisation use a fixed random seed of 42 to ensure reproducibility.
Input features are standardised using StandardScaler, transforming each feature to zero mean and unit variance. The scaler is fitted on the training set only and then applied to the validation and test sets to avoid information leakage.
@@ 4 Results
This section reports the performance of the supervised classifier in distinguishing between the four EOS families within the constructed synthetic dataset. The input features include the gravitational mass, stellar radius, fundamental $f$-mode frequency, mode quadrupole coefficient $Q$, redshift, damping time, and characteristic gravitational-wave strain. The output classes correspond to four equation-of-state (EOS) families: nucleonic matter, strange matter, dark-matter–admixed matter, and hyperonic matter.
@@ 4.2 Overall classification performance
The classifier achieves a test-set accuracy of $97.4\%$ in the baseline configuration, evaluated on a stratified held-out test set containing $N_{\text{test}}=116$ samples. This performance reflects separability within the controlled model space considered here.
The detailed class-wise performance is summarised by the confusion matrix shown in Fig. 3.
@@ 4.6 Feature importance
These results indicate that oscillation-related observables provide complementary discriminatory information beyond static mass–radius trends within the present modelling assumptions. Their enhanced importance reflects sensitivity to stellar compactness and internal structure, which vary across EOS families. However, the degree of separability necessarily depends on the specific EOS realisations adopted and the fixed transport prescriptions used in constructing the dataset.
@@ 5 Discussion and Conclusions
Because the oscillation quantities are computed under fixed transport assumptions inherited from Ref. Husain, (2026), the classification performance reported here should be interpreted within this consistent modelling framework.
Different matter compositions can give rise to very similar effective equations of state and macroscopic stellar properties (see e.g. Alford et al., (2005) for hybrid stars that can masquerade as normal stars), leading to intrinsic degeneracies in classification. The present results should therefore be interpreted as conditional on the assumed EOS families and modelling prescriptions. The classifier identifies regions of distinguishability within this controlled theoretical space but does not imply unique composition determination for arbitrary EOS constructions.
@@
This work provides a minimal and reproducible computational baseline for studying model separability in neutron-star physics. The results demonstrate that supervised machine learning can effectively map regimes of distinguishability and degeneracy across representative EOS families. Extensions incorporating broader EOS ensembles, varying transport prescriptions, or samples drawn from observational posteriors will be necessary to assess robustness under more general conditions.
Code and Data Availability
The datasets and the Python framework used for the MLP classification, along with the scripts for stellar structure configurations, are available at https://github.com/wasif553/ML-NeutronStar-Classification.